# Connecting Federated ADMM to Bayes

### v1

@@ 5 Conclusions and Future Work
In this paper, we provide new connections between two distinct and previously unrelated federated learning approaches based on ADMM and Variational Bayes (VB), respectively. Our key result shows that the dual variables in ADMM naturally emerge through the site parameters used in VB. We first show this for the isotropic Gaussian case and then extend it to multivariate Gaussians. The latter is used to derive a new variant of ADMM where learned covariances incorporate uncertainty and are used as preconditioners. We also derive a new functional regularisation extension. Numerical experiments validate the improvements obtained in performance.
@@
There are many avenues for future work. We believe this connection holds beyond federated learning setting, and we hope to explore this in the future. This could give rise to a new set of algorithms that combine duality and uncertainty to solve challenging problems in machine learning, optimisation, and other fields. There are also several extensions for the federated case. Our current experiments sample all clients in every round of communication, and future experiments could relax this. We expect that our methods will still work well, just like algorithmically-related algorithms such as FedDyn. Future work can also analyse convergence rates of FedLap, following similar theoretical assumptions to those for FedDyn and FedADMM. We could also explore differentially-private versions of the algorithm, following differentially-private versions of PVI (Heikkilä et al., 2023). We also expect that our method will perform well in continual federated learning, using Bayesian continual learning techniques (Kirkpatrick et al., 2017; Pan et al., 2020; Titsias et al., 2020). Lastly, future work can explore connections between Federated Distillation and FedLap-Func.
@@ Appendix B Derivation of FedLap-Cov Updates
$\mathbb{D}_{KL}[q\,\|\,q_{g}]=\mbox{$\frac{1}{2}$}\|\bm{m}-\bm{w}_{g}\|^{2}_{\bm{S}_{g}}+\mbox{Tr}\left(\bm{S}_{g}\bm{S}^{-1}\right)-\log|\bm{S}_{g}\bm{S}^{-1}|+\text{const.}$.
@@
$\begin{split}(\bm{w}_{k},\bm{S}_{k})&\leftarrow\arg\min_{\bm{m},\bm{S}}\;\ell_{k}(\bm{m})+\mbox{Tr}\left[\mathbf{H}_{k}(\bm{m}_{\text{old}})\bm{S}^{-1}\right]+\bm{v}_{k}^{\top}\bm{m}-\mbox{$\frac{1}{2}$}\mbox{Tr}\left[\bm{V}_{k}\left(\bm{m}\bm{m}^{\top}+\bm{S}^{-1}\right)\right]\\ &\qquad\qquad\qquad\qquad\qquad\qquad+\mbox{$\frac{1}{2}$}\|\bm{m}-\bm{w}_{g}\|^{2}_{\bm{S}_{g}}+\mbox{Tr}\left(\bm{S}_{g}\bm{S}^{-1}\right)-\log|\bm{S}_{g}\bm{S}^{-1}|.\end{split}$ (19)
The dual term above (the fourth term) is expanded using the following identity,
@@
$\begin{split}\bm{w}_{k}&\leftarrow\arg\min_{\bm{m}}\;\ell_{k}(\bm{m})+\bm{v}_{k}^{\top}\bm{m}-\mbox{$\frac{1}{2}$}\bm{m}^{\top}\bm{V}_{k}\bm{m}+\mbox{$\frac{1}{2}$}\|\bm{m}-\bm{w}_{g}\|^{2}_{\bm{S}_{g}}\\ \bm{S}_{k}&\leftarrow\arg\min_{\bm{S}}\;\mbox{$\frac{1}{2}$}\mbox{Tr}\left[\mathbf{H}_{k}(\bm{m}_{\text{old}})\bm{S}^{-1}\right]-\mbox{$\frac{1}{2}$}\mbox{Tr}\left[\bm{V}_{k}\bm{S}^{-1}\right]+\mbox{Tr}\left(\bm{S}_{g}\bm{S}^{-1}\right)-\log|\bm{S}_{g}\bm{S}^{-1}|.\end{split}$ (20)
@@
$\mathbf{H}_{k}(\bm{m}_{\text{old}})-\bm{V}_{k}+\bm{S}_{g}-\bm{S}_{k}=0\quad\implies\quad\bm{S}_{k}=\mathbf{H}_{k}(\bm{m}_{\text{old}})-\bm{V}_{k}+\bm{S}_{g}$
In the main update, we set $\bm{m}_{\text{old}}$ to be the most recent value of $\bm{m}_{k}$ before updating it.
@@ C.2 Deriving the FedLap-Func algorithm
We derive the FedLap-Func equations (Eq. 17) from Eqs. 15 and 16. We note that we can derive a combined version of FedLap-Cov and FedLap-Func, but for simplicity we focus on the non-covariance case.
@@
We include this function-space term in both the client update for $\bm{w}_{k}$, as well as the global server update for $\bm{w}_{g}$. Using Sec. C.1, we add and subtract the same terms, giving us,
@@
Eq. 25 is simpler than Eq. 16, as the two weight-space terms have the same $(\bm{w}-\bm{w}_{g})$ quadratic form. Therefore we can straightforwardly consider this as changing the global server precision $\bm{S}_{g}$ to remove the contribution from $i\in\mathcal{M}_{k},\forall k$. When we take the approximation that all $\bm{S}=\delta\mathbf{I}$, we again set the Hessians to all be 0, and this simplifies the equation further. Plugging this into the first client update in Eq. 9, and multiplying all function-space terms by a hyperparameter $\tau$, gives us our final FedLap-Func algorithm.
@@ Appendix D Benefits of FedLap over FedDyn and FedADMM
In this section, we show two simple experiments showcasing why FedLap has better properties than FedDyn and FedADMM. First, in Sec. D.1 we show on a homogoneous MNIST split that FedDyn requires an additional weight-decay term in the local loss during training in order for it to perform well, unlike FedLap. Second, in Sec. D.2 we show how FedLap targets a better global loss (with weight-decay incorporated already), and gets closer as the number of communication rounds increases, unlike FedDyn and FedADMM, which target a worse $\bm{w}_{g}^{*,\delta=0}$.
@@ D.1 Importance of local weight-decay in FedDyn
without weight-decay, the algorithm gets worse over communication rounds (after quickly reaching decent performance).
@@ D.2 FedLap targets a better global loss than FedDyn/FedADMM
In this section, we show a simple setting showcasing the potential benefit of FedLap over FedDyn and FedADMM. We take the UCI Credit dataset (a binary classification task), split the data equally (and homogeneously) into just two clients, and train a logistic regression classifier, ensuring we train until convergence both locally (full-batch, Adam learning rate 0.001, 1000 local epochs/iterations) and for a large number of communication rounds (1000 communication rounds). FedADMM is always slower and worse-performing than FedDyn, so we only consider FedDyn next. Note that in this section, we do not include weight-decay in the local optimisation of FedDyn, because FedDyn’s theory does not include this, and this setting is very simple.
First, we note that ideal global performance (of $\bm{w}_{g}^{*}$) is better with $\delta=1$ (giving a test nll of 0.337) than $\delta=0$ (test nll 0.420). Therefore the $\bm{w}_{g}^{*}$ that FedLap targets is already a better target than that of FedDyn (and FedADMM).
Secondly, FedLap is quicker to target its ideal global parameter than FedDyn. FedLap reaches within 1% of its ideal $\bm{w}^{*}$ train loss within just 2 communication rounds, and continues getting closer as we train for longer (up to the 1000 rounds). This is despite setting different values of $\delta$ (meaning that FedLap might be targeting worse values of $\bm{w}_{g}^{*}$).
On the other hand, how close FedDyn gets to its own (worse) ideal train nll depends on the value of its hyperparameter $\delta^{\text{FedADMM}}$: the best value takes 78 rounds to get to within 1% of its optimal train nll. Larger values of the hyperparameter do not reach within 1% within 1000 rounds: they lead to quick improvement earlier in training, but then performance (and train nll) gets worse over time.
@@ E.2 UCI Credit
We set local Adam learning rate at $10^{-3}$, and minibatch size to 4 (ensuring there is gradient noise always). All methods have a hyperparameter sweep over number of local epochs = $[5,10,20]$. We perform a sweep over the $\delta$ hyperparameter for FedProx, FedLap, FedLap-Cov and FedLap-Func over $\delta=[10,1,10^{-1}]$. For FedDyn the sweep is over $\delta=[10^{-3},10^{-4},10^{-5}]$, and local weight-decay sweep is over $[10^{-4},10^{-5},10^{-6}]$. For FedLap and FedLap-Func, we set damping factor $\rho=N_{k}/N$, and for FedLap-Cov, $\rho=1/K$. For FedLap-Func, we set $\tau=1$ , and run global server optimisation for 5000 iterations with a learning rate of 0.001. Each run took up to 4 minutes on a standard laptop CPU.
@@ E.3 FLamby-Heart
We set batch size to 4 (like in du Terrail et al. [2022]), but use the Adam optimiser locally. we perform hyperparameter sweeps over number of local epochs = $[1,5,10]$, and (Adam) local learning rate = $[10^{-3},10^{-2}]$. For FedProx, FedLap, and FedLap-Cov, we perform a sweep over $\delta=[100,10,1,10^{-1},10^{-2}]$. For FedDyn, this is over $\delta=[10,1,10^{-1},10^{-2},10^{-3}]$, and FedDyn’s local weight-decay sweep is over $[0,10^{-2},10^{-3},10^{-4}]$. For FedLap, we set damping factor $\rho=N_{k}/N$, and for FedLap-Cov, $\rho=1/K$. Each run took less than 1 minute on a standard laptop CPU.
@@ E.4 MNIST
We use the full MNIST dataset (the model gets 98.3% accuracy).
Homogoneous and heterogeneous MNIST. Batch size is set to 32. We train a two-hidden layer multi-layer perceptron, with 200 hidden units in the first layer, and 100 units in the second layer.
For all methods, we perform hyperparameter sweeps over number of local epochs = $[1,5,10]$, and (Adam) local learning rate = $[10^{-3},10^{-2}]$. For FedProx, FedLap, FedLap-Cov and FedLap-Func, we perform a sweep over $\delta=[1,10^{-1},10^{-2}]$. For FedDyn, this is over $\delta=[10^{-4},10^{-5},10^{-6}]$, and FedDyn’s local weight-decay sweep is over $[10^{-4},10^{-5},10^{-6}]$. For FedLap and FedLap-Func, we set damping factor $\rho=N_{k}/N$, and for FedLap-Cov, $\rho=1/K$. For FedLap-Func, we sweep over $\tau_{f}=[1,10^{-1},10^{-2}]$ , and run global server optimisation for 3000 iterations with a learning rate of 0.0005. In the first round of communication, each client randomly picks 1 point per class present in the client, and transmits the input (not true label) to the server. Each run took up to 2 hours on a standard laptop CPU.
@@ E.5 Fashion MNIST
We use a random $10\%$ of the data (different subset in every random seed) in FMNIST to simulate having less data (the model gets 85% accuracy).
Homogoneous and heterogeneous Fashion MNIST. Batch size is set to 32. We train a two-hidden layer multi-layer perceptron, with 200 hidden units in the first layer, and 100 units in the second layer. We use a random 10% of Fashion MNIST every random seed, and average across (the same) three random seeds. Local Adam learning rate is set to $10^{-3}$.
For all methods, we perform hyperparameter sweeps over number of local epochs = $[1,5,10]$. For FedProx, FedLap, FedLap-Cov and FedLap-Func, we perform a sweep over $\delta=[1,10^{-1},10^{-2},10^{-3},10^{-4}]$. For FedDyn, this is over $\delta=[10^{-2},10^{-3},10^{-4},10^{-5},10^{-6}]$, and FedDyn’s local weight-decay sweep is over $[10^{-4},10^{-5},10^{-6}]$. For FedLap and FedLap-Func, we set damping factor $\rho=N_{k}/N$, and for FedLap-Cov, $\rho=1/K$. For FedLap-Func, we sweep over $\tau_{f}=[1,10^{-1},10^{-2},10^{-3},10^{-4}]$ , and run global server optimisation for 5000 iterations with a learning rate of 0.001. In the first round of communication, each client randomly picks 1 point per class present in the client, and transmits the input (not true label) to the server. Each run took up to 1 hour 20 minutes on a standard laptop CPU.
@@ E.6 CIFAR-10
For all methods, we perform hyperparameter sweeps over number of local epochs = $[5,10,20]$. For FedProx, we perform a sweep over $\delta=[1,10^{-1},10^{-2},10^{-3}]$. For FedLap, FedLap-Cov and FedLap-Func, we perform a sweep over $\delta=[1,10^{-1},10^{-2}]$. For FedDyn, this is over $\delta=[10^{-3},10^{-4},10^{-5}]$, and FedDyn’s local weight-decay sweep is over $[10^{-4},10^{-5},10^{-6}]$.
For FedLap and FedLap-Func, damping $\rho$ was $N_{k}/N$ for first 10 rounds of training, then $1/K$ after that, and gradient clipping was used to stabilise training. For FedLap-Func, we sweep over $\tau_{f}=[10^{-2},10^{-4},10^{-6}]$ , and run global server optimisation for 500 iterations with a learning rate of 0.0005. In the first round of communication, each client randomly picks 2 points per class present in the client, and transmits the input (not true label) to the server. Each run took up to 10 hours on an A100 GPU.

### v2

@@ 5 Conclusions and Future Work
In this paper, we provide new connections between two distinct and previously unrelated federated learning approaches based on ADMM and Variational Bayes (VB), respectively. Our key result shows that the dual variables in ADMM naturally emerge through the site parameters used in VB. We first show this for the isotropic Gaussian case and then extend it to multivariate Gaussians. The latter is used to derive a new variant of ADMM where learned covariances incorporate uncertainty and are used as preconditioners. We also derive a new functional regularisation extension. Numerical experiments show that these improve performance.
@@
There are many avenues for future work. We believe this connection holds beyond the federated learning setting, and we hope to explore this in the future. This could give rise to a new set of algorithms that combine duality and uncertainty to solve challenging problems in machine learning, optimisation, and other fields. There are also several extensions for the federated case. Our current experiments update all clients in every round of communication, and future experiments could relax this. We expect that our methods will still work well, just like algorithmically-related algorithms such as FedDyn. Future work can also analyse convergence rates of FedLap, following similar theoretical assumptions to those for FedDyn and FedADMM. We could also explore differentially-private versions of the algorithm, following differentially-private versions of PVI (Heikkilä et al., 2023). We also expect that our method will perform well in continual federated learning, using Bayesian continual learning techniques (Kirkpatrick et al., 2017; Pan et al., 2020; Titsias et al., 2020). Lastly, future work can explore connections between Federated Distillation and FedLap-Func.
@@ Acknowledgments
Author Contributions Statement
List of authors: Siddharth Swaroop (SS), Mohammad Emtiyaz Khan (MEK), Finale Doshi-Velez (FDV).
SS led the project and conceived the original ideas with regular feedback from FDV and MEK. The emergence of dual variables through site parameter was hypothesised by MEK, and confirmed by SS. SS coded and ran all experiments. SS wrote a first draft of the paper with feedback from FDV, which was substantially revised by MEK with feedback from SS and FDV.
@@ Appendix B Derivation of FedLap-Cov Updates
$\mathbb{D}_{KL}[q\,\|\,q_{g}]=\mbox{$\frac{1}{2}$}\|\bm{m}-\bm{w}_{g}\|^{2}_{\bm{S}_{g}}+\mbox{$\frac{1}{2}$}\mbox{Tr}\left(\bm{S}_{g}\bm{S}^{-1}\right)-\mbox{$\frac{1}{2}$}\log|\bm{S}_{g}\bm{S}^{-1}|+\text{const.}$.
@@
$\begin{split}(\bm{w}_{k},\bm{S}_{k})&\leftarrow\arg\min_{\bm{m},\bm{S}}\;\ell_{k}(\bm{m})+\mbox{$\frac{1}{2}$}\mbox{Tr}\left[\mathbf{H}_{k}(\bm{m}_{\text{old}})\bm{S}^{-1}\right]+\bm{v}_{k}^{\top}\bm{m}-\mbox{$\frac{1}{2}$}\mbox{Tr}\left[\bm{V}_{k}\left(\bm{m}\bm{m}^{\top}+\bm{S}^{-1}\right)\right]\\ &\qquad\qquad\qquad\qquad\qquad+\mbox{$\frac{1}{2}$}\|\bm{m}-\bm{w}_{g}\|^{2}_{\bm{S}_{g}}+\mbox{$\frac{1}{2}$}\mbox{Tr}\left(\bm{S}_{g}\bm{S}^{-1}\right)-\mbox{$\frac{1}{2}$}\log|\bm{S}_{g}\bm{S}^{-1}|.\end{split}$ (19)
The $\bm{V}_{k}$ dual term above (the fourth term) is derived using the following identity,
@@
$\begin{split}\bm{w}_{k}&\leftarrow\arg\min_{\bm{m}}\;\ell_{k}(\bm{m})+\bm{v}_{k}^{\top}\bm{m}-\mbox{$\frac{1}{2}$}\bm{m}^{\top}\bm{V}_{k}\bm{m}+\mbox{$\frac{1}{2}$}\|\bm{m}-\bm{w}_{g}\|^{2}_{\bm{S}_{g}},\\ \bm{S}_{k}&\leftarrow\arg\min_{\bm{S}}\;\mbox{$\frac{1}{2}$}\mbox{Tr}\left[\mathbf{H}_{k}(\bm{m}_{\text{old}})\bm{S}^{-1}\right]-\mbox{$\frac{1}{2}$}\mbox{Tr}\left[\bm{V}_{k}\bm{S}^{-1}\right]+\mbox{$\frac{1}{2}$}\mbox{Tr}\left(\bm{S}_{g}\bm{S}^{-1}\right)-\mbox{$\frac{1}{2}$}\log|\bm{S}_{g}\bm{S}^{-1}|.\end{split}$ (20)
@@
$\mathbf{H}_{k}(\bm{m}_{\text{old}})-\bm{V}_{k}+\bm{S}_{g}-\bm{S}_{k}=0\quad\implies\quad\bm{S}_{k}=\mathbf{H}_{k}(\bm{m}_{\text{old}})-\bm{V}_{k}+\bm{S}_{g}.$
In the main update, we set $\bm{m}_{\text{old}}$ to be the most recent value of $\bm{m}_{k}$ before updating it. This ensures that the Hessian is estimated at the most recent parameters and so is the best estimate possible.
@@ C.2 Deriving the FedLap-Func algorithm
We derive the FedLap-Func equations (Eq. 17) from Eqs. 15 and 16. We note that we can derive a combined version of FedLap-Cov and FedLap-Func, but for simplicity we will focus on the non-covariance case.
@@
We include this function-space term in both the client update for $\bm{w}_{k}$ and the global server update for $\bm{w}_{g}$. We illustrate this with the global server update. Using Sec. C.1 gives us,
@@
Eq. 25 is simpler than Eq. 16, as the two weight-space terms have the same $(\bm{w}-\bm{w}_{g})$ quadratic form. Therefore we can straightforwardly consider this as changing the global server precision $\bm{S}_{g}$ to remove the contribution from $i\in\mathcal{M}_{k},\forall k$. When we take the approximation that all $\bm{S}=\delta\mathbf{I}$, we again set the Hessians to all be 00, and this simplifies the equation further. Plugging this into the first client update in Eq. 9, and multiplying all function-space terms by a hyperparameter $\tau$, gives us our final FedLap-Func algorithm.
@@ Appendix D Benefits of FedLap over FedDyn and FedADMM
In this section, we show two simple experiments showcasing why FedLap has better properties than FedDyn and FedADMM. First, in Sec. D.1 we show on a homogoneous MNIST split that FedDyn requires an additional weight-decay term in the local loss during training in order for it to perform well, unlike FedLap. Second, in Sec. D.2 we show how FedLap targets a better global loss (with weight-decay incorporated already), and gets closer as the number of communication rounds increases, unlike FedDyn and FedADMM, which target a worse global optimum (with no weight-decay in the global loss).
@@ D.2 FedLap targets a better global loss than FedDyn/FedADMM
In this section, we present a simple setting showcasing the potential benefits of FedLap over FedDyn (and FedADMM). We take the UCI Credit dataset (a binary classification task), split the data equally (and homogeneously) into just two clients, and train a logistic regression classifier, ensuring we train until convergence both locally (full-batch, Adam learning rate 0.001, 1000 local epochs/iterations) and for a large number of communication rounds (1000 communication rounds). FedADMM is always slower and worse-performing than FedDyn, so we only consider FedDyn. Note that in this section, we do not include weight-decay in the local optimisation of FedDyn, because FedDyn’s theory does not include this, and this setting is very simple.
First, we note that ideal global performance (of $\bm{w}_{g}^{*}$) is better with $\delta=1$ (giving a test nll of 0.337) than $\delta=0$ (test nll 0.420). Therefore the $\bm{w}_{g}^{*}$ that FedLap targets is already a better target than that of FedDyn (and FedADMM), which targets a global optimum with $\delta=0$.
Secondly, FedLap is quicker to target its ideal global parameter than FedDyn. FedLap reaches within 1% of its ideal $\bm{w}^{*}$ train loss within just 2 communication rounds, and continues getting closer as we train for longer (up to the 1000 rounds). This happens even when setting different values of $\delta$ (meaning that FedLap might be targeting unideal values of $\bm{w}_{g}^{*}$).
On the other hand, how close FedDyn gets to its own (worse) ideal train nll depends on the value of its hyperparameter $\alpha$: the best value takes 78 rounds to get to within 1% of its optimal train nll. Larger values of the hyperparameter do not reach within 1% within 1000 rounds: they lead to quick improvement earlier in training, but then performance (and train nll) gets worse over time.
@@ E.2 UCI Credit
We set local Adam learning rate at $10^{-3}$, and minibatch size to 4 (ensuring there is gradient noise always). All methods have a hyperparameter sweep over number of local epochs = $[5,10,20]$. We perform a sweep over the $\alpha$ or $\delta$ hyperparameter for FedProx, FedLap, FedLap-Cov and FedLap-Func over $[10,1,10^{-1}]$. For FedDyn the sweep is over $\alpha=[10^{-3},10^{-4},10^{-5}]$, and local weight-decay sweep is over $[10^{-4},10^{-5},10^{-6}]$. For FedLap and FedLap-Func, we set damping factor $\rho=N_{k}/N$, and for FedLap-Cov, $\rho=1/K$. For FedLap-Func, we set $\tau=1$ , and run global server optimisation for 5000 iterations with a learning rate of 0.001. Each run took up to 4 minutes on a standard laptop CPU.
@@ E.3 FLamby-Heart
We set batch size to 4 (like in du Terrail et al. [2022]), but use the Adam optimiser locally. We perform hyperparameter sweeps over number of local epochs = $[1,5,10]$, and (Adam) local learning rate = $[10^{-3},10^{-2}]$. For FedProx, FedLap, and FedLap-Cov, we perform a sweep over $\delta\,(\text{or }\alpha)=[100,10,1,10^{-1},10^{-2}]$. For FedDyn, this is over $\alpha=[10,1,10^{-1},10^{-2},10^{-3}]$, and FedDyn’s local weight-decay sweep is over $[0,10^{-2},10^{-3},10^{-4}]$. For FedLap, we set damping factor $\rho=N_{k}/N$, and for FedLap-Cov, $\rho=1/K$. Each run took less than 1 minute on a standard laptop CPU.
@@ E.4 MNIST
We use the full MNIST dataset. Batch size is set to 32. We train a two-hidden layer multi-layer perceptron, with 200 hidden units in the first layer, and 100 units in the second layer (the model gets 98.3% accuracy on all data).
For all methods, we perform hyperparameter sweeps over number of local epochs = $[1,5,10]$, and (Adam) local learning rate = $[10^{-3},10^{-2}]$. For FedProx, FedLap, FedLap-Cov and FedLap-Func, we perform a sweep over $\delta\,(\text{or }\alpha)=[1,10^{-1},10^{-2}]$. For FedDyn, this is over $\alpha=[10^{-4},10^{-5},10^{-6}]$, and FedDyn’s local weight-decay sweep is over $[10^{-4},10^{-5},10^{-6}]$. For FedLap and FedLap-Func, we set damping factor $\rho=N_{k}/N$, and for FedLap-Cov, $\rho=1/K$. For FedLap-Func, we sweep over $\tau_{f}=[1,10^{-1},10^{-2}]$ , and run global server optimisation for 3000 iterations with a learning rate of 0.0005. Each run took up to 2 hours on a standard laptop CPU.
@@ E.5 Fashion MNIST
We use a random 10% of Fashion MNIST every random seed, and average across (the same) three random seeds. Batch size is set to 32. We train a two-hidden layer multi-layer perceptron, with 200 hidden units in the first layer, and 100 units in the second layer (the model gets 85% accuracy). Local Adam learning rate is set to $10^{-3}$.
For all methods, we perform hyperparameter sweeps over number of local epochs = $[1,5,10]$. For FedProx, FedLap, FedLap-Cov and FedLap-Func, we perform a sweep over $\delta\,(\text{or }\alpha)=[1,10^{-1},10^{-2},10^{-3},10^{-4}]$. For FedDyn, this is over $\alpha=[10^{-2},10^{-3},10^{-4},10^{-5},10^{-6}]$, and FedDyn’s local weight-decay sweep is over $[10^{-4},10^{-5},10^{-6}]$. For FedLap and FedLap-Func, we set damping factor $\rho=N_{k}/N$, and for FedLap-Cov, $\rho=1/K$. For FedLap-Func, we sweep over $\tau_{f}=[1,10^{-1},10^{-2},10^{-3},10^{-4}]$ , and run global server optimisation for 5000 iterations with a learning rate of 0.001. Each run took up to 1 hour 20 minutes on a standard laptop CPU.
@@ E.6 CIFAR-10
For all methods, we perform hyperparameter sweeps over number of local epochs = $[5,10,20]$. For FedProx, we perform a sweep over $\alpha=[1,10^{-1},10^{-2},10^{-3}]$. For FedLap, FedLap-Cov and FedLap-Func, we perform a sweep over $\delta=[1,10^{-1},10^{-2}]$. For FedDyn, this is over $\alpha=[10^{-3},10^{-4},10^{-5}]$, and FedDyn’s local weight-decay sweep is over $[10^{-4},10^{-5},10^{-6}]$.
For FedLap and FedLap-Func, damping $\rho$ was $N_{k}/N$ for first 10 rounds of training, then $1/K$ after that, and gradient clipping was used to stabilise training. For FedLap-Func, we sweep over $\tau_{f}=[10^{-2},10^{-4},10^{-6}]$ , and run global server optimisation for 500 iterations with learning rate 0.0005. We assume that each client has 2 randomly-selected points per class shared with the global server (all other dataset/benchmark settings have 1 point per class per client). Each run took up to 10 hours on an A100 GPU.